Dr Johnson invests $9,000 at the rate of 6% annual interest compounded semi annually for 4 years. what is the future value
Answer #1

Rate = 6% / 2 = 3%

Number of periods = 4 * 2 = 8

future value = Present value (1 + r)n

future value = 9000 (1 + 0.03)8

future value = 9000 * 1.26677

future value = $11,400.93